### v1

# CARFT: Boosting LLM Reasoning via Contrastive Learning with Annotated Chain-of-Thought-based Reinforced Fine-Tuning

Abstract
Reasoning capability plays a significantly critical role in the the broad applications of Large Language Models (LLMs). To enhance the reasoning performance of LLMs, diverse Reinforcement Learning (RL)-based fine-tuning approaches have been proposed to address the limited generalization capability of LLMs trained solely via Supervised Fine-Tuning (SFT). Despite their effectiveness, two major limitations hinder the advancement of LLMs. First, vanilla RL-based approaches ignore annotated Chain-of-Thought (CoT) and incorporate unstable reasoning path sampling, which typically results in model collapse, unstable training process, and suboptimal performance. Second, existing SFT approaches generally overemphasize the annotated CoT, potentially leading to performance degradation due to insufficient exploitation of potential CoT. In this paper, we propose a Contrastive learning with annotated CoT-based Reinforced Fine-Tuning approach, i.e., CARFT, to enhance the reasoning performance of LLMs while addressing the aforementioned limitations. Specifically, we propose learning a representation for each CoT. Based on this representation, we design novel contrastive signals to guide the fine-tuning process. Our approach not only fully exploits the available annotated CoT but also stabilizes the fine-tuning procedure by incorporating an additional unsupervised learning signal. We conduct comprehensive experiments and in-depth analysis with three baseline approaches, two foundation models, and two datasets to demonstrate significant advantages of CARFT in terms of robustness, performance (up to 10.15%), and efficiency (up to 30.62%). Code is available at https://github.com/WNQzhu/CARFT.
CARFT: Boosting LLM Reasoning via Contrastive Learning with Annotated Chain-of-Thought-based Reinforced Fine-Tuning
Wenqiao Zhu1, 2, $\dagger$, Ji Liu1, Rongjuncheng Zhang1, Haipang Wu1, Yulun Zhang2 1HiThink Research, 2Shanghai Jiao Tong University
[…]
1 Introduction
The reasoning capability of Large Language Models (LLMs) stands as a critical component, driving an extensive array of potential applications, which span mathematical problem (Wang et al., 2024; Luo et al., 2023), financial analysis (Yang et al., 2023; Zhang et al., 2023a), and medical applications (Singhal et al., 2022), etc. The advent of reasoning LLMs, e.g., OpenAI o1 (Jaech et al., 2024), OpenAI o3 (OpenAI, 2025), Llama-Nemotron (Bercovich et al., 2025), Claude 3.7 (Anthropic, 2024), and DeepSeek R1 (DeepSeek-AI, 2025), has significantly heightened the interest in exploring the reasoning capabilities of LLMs across both academic and industrial sectors. Additionally, given the straightforward verification of answers, the task of solving mathematical problems has emerged as a pivotal domain in the study of LLM reasoning capacities.
One of the conventional strategies for augmenting the reasoning capabilities of LLMs is Supervised Fine-Tuning (SFT). SFT entails fine-tuning LLMs with training samples that incorporate annotated Chain-of-Thought (CoT) (Wei et al., 2022). In a training dataset $\mathcal{D}_{\text{train}}$, each training sample is structured as a tuple $(\mathbf{x},\mathbf{c},\mathbf{y})$, where $\mathbf{x}$ represents the input question, $\mathbf{c}$ represents the annotated CoT, and $\mathbf{y}$ denotes the correct ground truth answer. CoT $\mathbf{c}$ in the training sample is generally written or labeled by experienced experts or high-end LLMs, which is highly valuable for the fine-tuning of LLMs.
SFT-based reasoning enhancement approaches only exploit a single annotated CoT for each question within the training dataset. However, multiple CoTs (Zhang et al., 2023b) exist for each question. Hence, conducting SFT with only a single annotated CoT in the training dataset may limit the generalization capability of LLMs.
Method A-CoT SG-CoT
[…]
Table 1: An overview of whether methods employ Annotated-CoT (A-CoT) or Self-Generated CoT (SG-CoT).
To address the limitations of SFT-based methods, Reinforcement Learning (RL)-based fine-tuning approaches emerge (Luong et al., 2024; Shao et al., 2024; Liu et al., 2025). A prominent and state-of-the-art RL-based fine-tuning approach is ReFT (Luong et al., 2024), which incorporates online RL approach, i.e., Proximal Policy Optimization (PPO) (Schulman et al., 2017), to dynamically sample CoTs at each training step. This mechanism enables ReFT to leverage multiple CoTs, thereby improving the generalization capability of LLMs.
Figure 1: Comparison between SFT, ReFT, and CARFT on the exploration of CoT.
Despite their effectiveness, two major limitations still exist with the existing RL-based fine-tuning approaches, which hinders the advancement of LLMs. First, existing RL-based approaches solely rely on on-policy sampled CoTs without considering the highly valued annotated CoTs while enhancing reasoning performance. Due to the reward hacking problem (Skalse et al., 2022), such sampled CoTs may not be valid or correct, potentially degrading model performance. Second, existing approaches suffer from unstable training. The inherent exploratory nature of RL can lead to model collapse, i.e., a phenomenon where the behavior of LLMs significantly deteriorates during training. This instability can severely impact the performance of LLMs and result in undesirable outputs.
To address these challenges, we propose a novel Contrastive learning with Annotated CoT-based Reinforced Fine-Tuning approach, i.e., CARFT, which effectively leverages the valuable annotated CoTs in the training dataset while sampling other potential CoTs so as to achieve superb performance. CARFT begins with learning a unified representation for each CoT, encompassing both high-quality annotated CoTs and on-policy sampled CoTs. Based on this representation, we design contrastive signals to improve both the reasoning performance and the stability of the fine-tuning process. Specifically, we propose exploiting a masked loss, e.g., InfoNCE (Chen et al., 2020), to utilize the unified representation to generate the contrastive signal. This contrastive signal serves as a guiding mechanism for the on-policy CoT sampling process, helping to stabilize the fine-tuning of LLMs while maximizing the utilization of information from the annotated CoT. Table 1 and Figure 1 illustrate the working characteristic of CARFT. In addition, we propose embedding-enhanced partial reward to further improve the performance. The key contributions of this paper are summarized as follows:
  • We propose a novel contrastive learning-based framework with an original contrastive signal construction method that fully exploits annotated CoTs to improve both the performance and the stability in the fine-tuning of LLMs.
  • We design an embedding-enhanced partial reward so as to further improve the stability in the reinforced fine-tuning process and to achieve superb performance of LLMs.
  • We conduct extensive experiments and thorough ablation studies to demonstrate the effectiveness of CARFT compared with three baseline approaches, two foundation models, and two datasets. Extensive experimental results demonstrate that CARFT significantly outperformances baselines in terms of effectiveness (up to 10.15%) and robustness.
2 Related Work
Recent years have witnessed widespread application in Natural Language Processing (NLP), particularly in the domains of preference optimization (Stiennon et al., 2020; Rafailov et al., 2023; Gheshlaghi Azar et al., 2024; Zhu et al., 2025) and reasoning (Luong et al., 2024; Shao et al., 2024; Yu et al., 2025; Liu et al., 2025). These methods typically follow a standard three-stage pipeline: (1) SFT, (2) reward modeling, and (3) RL-based optimization. A key distinction among these approaches lies in how the reward signal is obtained. In preference optimization, reward models are learned from human feedback, while in mathematical reasoning tasks, rule-based methods are typically exploited to construct reward signals, as ground-truth answers can be explicitly verified. Within the context of preference optimization, Direct Preference Optimization (DPO) (Rafailov et al., 2023) has emerged as an effective algorithm that avoids the need for explicit reward model training. However, due to its offline nature (Feng et al., 2024), DPO may struggle to explore diverse CoTs (Luong et al., 2024). As a result, on-policy approaches, e.g., GRPO Shao et al. (2024), DAPO (Yu et al., 2025), Dr.GRPO (Liu et al., 2025), and ReFT Luong et al. (2024), are generally employed to better explore such diversity in reasoning.
On-policy approaches utilize multiple rollouts to estimate the Generalized Advantage Estimation (GAE). DAPO and Dr.GRPO are both improved variants of GRPO. Specifically, DAPO is designed for long-CoT scenarios and introduces four key techniques: higher clipping, dynamic sampling, token-level policy gradient loss, and overlong reward shaping. On the other hand, Dr.GRPO improves upon GRPO by eliminating the bias present in the original method. While these approaches are effective, they come with the trade-off of increased computational complexity. In contrast, ReFT (Luong et al., 2024) utilizes only a single on-policy sample per step, making it significantly computationally efficient.
Despite their strengths, these approaches rely solely on on-policy sampling, ignoring potentially valuable annotated CoTs already present in the training data. Moreover, model collapse occurs frequently within the reinforced fine-tuning process with the existing approaches.
3 Method
In this section, we first present the preliminary of reinforced LLM fine-tuning. Then, we detail CARFT, including the contrastive learning-based framework with an original contrastive signal construction method and an embedding enhanced partial reward method.
Figure 2: The framework CARFT is composed of two sequential stages: (i) supervised fine-tuning (SFT), followed by (ii) contrastive feedback.
3.1 Preliminary of Reinforced Fine-Tuning
Reinforced fine-tuning incorporates feedback signals derived from either a learned reward model or predefined rules to guide the training of LLMs. Given an input prompt $x$ and the corresponding response $y$ produced by a LLM, the objective is to maximize the expected cumulative reward, which is formally formulated as:
$\max_{\pi_{\theta}}$ $\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)\right]$
$-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right],$ (1)
where $r(x,y)$ denotes the reward function, $\theta$ represents the parameters of the LLM, $\pi_{\theta}$ refers to the learnable target policy, $\pi_{\text{ref}}$ corresponds to a reference policy, i.e., typically the initially pre-trained LLM, exploited to stabilize training, and $\beta$ denotes the coefficient of KL-divergence, which encourages the updated policy to stay close to the original distribution.
A commonly used algorithm in this setting is PPO (Schulman et al., 2017), which employs GAE (Schulman et al., 2015) for stable gradient updates. In particular, the advantage at time step $t$ is calculated as:
$\hat{A}_{t}=\sum_{l=0}^{L-t}(\gamma\lambda)^{l}\sigma_{t+l},$ (2)
where $L$ represents the maximum length of token sequence, $\sigma_{t^{\prime}}$ is the Temporal Difference (TD) residual at timestep $t^{\prime}$, $\lambda\in(0,1]$ serves as the GAE discount factor, and $\gamma\in[0,1]$ controls the discounting of TD residuals. $\sigma_{t^{\prime}}$ is formulated as:
$\sigma_{t^{\prime}}=-V_{\phi}(s_{t^{\prime}})+r_{\text{total}}(s_{t^{\prime}},a_{t^{\prime}},s_{t^{\prime}+1})+\gamma V_{\phi}(s_{t^{\prime}}),$
where $s_{t^{\prime}}$ denotes the state at time step $t^{\prime}$, and $a_{t^{\prime}}$ represents the action at time step $t^{\prime}$, $\gamma$ is similar as defined in Formula 2, $r_{\text{total}}(\cdot)$ calculates the total reward, and $V_{\phi}(\cdot)$ estimates the state value from a given state with $\phi$ referring to a vector of policy parameters. Actions correspond to individual tokens selected from the vocabulary.
The total reward at the token level includes both the external reward signal and an internal regularization based on the Kullback–Leibler (KL) divergence between the current and reference policies, as defined as follows:
$r_{\text{total}}(s_{t^{\prime}},a_{t^{\prime}},s_{t^{\prime}+1})$ $=r(s_{t^{\prime}},a_{t^{\prime}})$
$+\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(\cdot|s_{t^{\prime}})||\pi_{\text{ref}}(\cdot|s_{t^{\prime}})\right],$
where $\pi_{\theta}$ represents the sampling actions policy with policy parameters $\phi$, and $\pi_{\text{ref}}$ corresponds to the sampling actions with a reference policy. Given GAE $\hat{A}_{t}$ and state value $V_{\phi}(s_{t})$, we can estimate the reward as defined in Formula 3:
$\hat{R}_{t}=\hat{A}_{t}+V_{\phi}(s_{t}).$ (3)
Under the PPO framework, the policy and value loss functions are separately formulated to ensure a stable and effective fine-tuning process. The policy loss is defined in Formula 3.1.
[…]
$\left.\left.\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi^{\text{old}}_{\theta}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right],$ (4)
[…]
$\left\|\left.\left.\text{clip}\left(\hat{R}_{t}-V_{\phi}(s_{t}),\hat{A}_{t}-\epsilon,\hat{A}_{t}+\epsilon\right)\right\|^{2}\right)\right],$
where $\epsilon$ is similar as defined in 3.1. Finally, the overall reinforcement learning loss combines both the policy loss and the value loss as defined in Formula 5:
$\mathcal{L}_{\text{RL}}=\mathcal{L}_{\text{policy}}+\alpha\mathcal{L}_{\text{value}},$ (5)
where $\alpha$ balances the relative importance of the policy and value losses within the reinforced fine-tuning process.
3.2 CARFT
While existing approaches either overemphasizes the annotated CoT (for SFT) or face challenges in achieving stable reinforced fine-tuning while ignoring annotated CoT (for existing RL-based approaches), we propose a novel contrastive learning-based approach, i.e., CARFT, to properly levarage the annotated CoTs so as to address this issue. In this section, we first present the overall workflow of CARFT. Then, we explain the CoT embeddings. Afterward, we propose masked contrastive signal construction approach in CARFT. Finally, we explain a novel embedding enhanced partial reward method.
3.2.1 Workflow
As shown in Figure 2, the overall workflow of CARFT consists of two sequential stages: the SFT stage and the reinforced fine-tuning stage.
SFT
We assume that each training sample in the training dataset is a triplet $(\mathbf{x},\mathbf{c},\mathbf{y})$, where $\mathbf{x}$ denotes the input question, $\mathbf{c}$ represents the annotated CoT, and $\mathbf{y}$ is the ground-truth answer. We carry out SFT with a few epochs to improve the instruction-following ability of the LLM.
Contrastive Feedback
Let $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ denote two distinct CoTs corresponding to the same input question $\mathbf{x}_{1}$, derived either from training examples or on-policy sampling, i.e., $(\mathbf{x}_{1},\mathbf{c}_{1},\mathbf{y}_{1})$ and $(\mathbf{x}_{1},\mathbf{c}_{2},\mathbf{y}_{1})$. We assume that $\mathbf{y}_{1}$ is the valid answer to $\mathbf{x}_{1}$. Since $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ pertain to the same input $\mathbf{x}_{1}$, we posit the existence of a conditional distribution $p_{1}(\mathbf{c}\mid\mathbf{h}_{1})$ such that both $\mathbf{c}_{1}\sim p_{1}(\mathbf{c}\mid\mathbf{h}_{1})$ and $\mathbf{c}_{2}\sim p_{1}(\mathbf{c}\mid\mathbf{h}_{1})$, where $\mathbf{h}_{1}$ denotes a latent variable associated with $\mathbf{x}_{1}$. Given that $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ are sampled from the same distribution, there should exist a similarity metric $m(\cdot,\cdot)$ under which the distance between $\mathbf{c}_{1}$ and $\mathbf{c}_{2}$ is smaller than the distance between $\mathbf{c}_{1}$ and any $\mathbf{c}_{i}$ drawn from a different distribution $p_{i,i\neq 1}(\mathbf{c}\mid\mathbf{h}_{i})$, with a high probability:
[…]
This insight provides two key advantages when incorporated as an unsupervised signal. First, it enables us to exploit the annotated CoTs in the training data in the reinforced fine-tuning process of LLMs. Second, it offers a guiding signal for CoT generation, helping to stabilize the reinforced fine-tuning process and to mitigate the risk of model collapse.
3.2.2 Chain-of-Thought Embeddings
Given a CoT $\mathbf{c}$ of length $L$, represented as:
$\mathbf{c}=\left[a_{1},a_{2},\cdots,a_{L}\right],$
we denote the corresponding token embeddings and state values as:
$\mathbf{H}=\left[H_{1},H_{2},\cdots,H_{L}\right]$
and
$\mathbf{V}_{\phi}=\left[V_{\phi}(1),V_{\phi}(2),\cdots,V_{\phi}(L)\right],$
respectively.
To obtain a compact representation of the entire CoT, we compute a weighted sum of the token embeddings using the softmax-normalized state values, as defined in Formula 6.
$\mathbf{e}=\sum\text{Softmax}(\mathbf{V}_{\phi})\odot\mathbf{H},$ (6)
where $\odot$ denotes element-wise multiplication between the state values and the corresponding embedding vectors.
In practice, in order to reduce memory consumption, we first project each embedding $H_{i}$ into a lower-dimensional space using a simple single-layer MultiLayer Perceptron (MLP), denoted by $\text{proj}(\cdot)$. The projected embeddings are then exploited in place of the original ones:
$\mathbf{H}=\left[\text{proj}(H_{1}),\text{proj}(H_{2}),\cdots,\text{proj}(H_{L})\right].$
3.2.3 Masked Contrastive Signal Construction
In this section, we design two types of contrastive signals for reinforced fine-tuning, i.e., positive and negative. We denote the signal related to CoT that results in a correct answer by positive signal, and that results in a wrong answer by negative signal.
Positive Signal
Given a batch of training samples $\{\mathbf{x}_{i},\mathbf{c}_{i}^{\text{annotated}},\mathbf{y}_{i}\}_{1}^{B}$ with $B$ presenting the batch size, we conduct LLM self-generation to generate a batch of rollout CoTs, i.e., $\{\mathbf{x}_{i},\mathbf{c}_{i}^{\text{rollout}},\mathbf{y}_{i}\}_{1}^{B}$. By employing the CoT embedding module, we could get embeddings of the annotated CoTs $\{\mathbf{e}^{\text{annotated}}_{i}\}_{1}^{B}$ and rollout CoT $\{\mathbf{e}^{\text{rollout}}_{i}\}_{1}^{B}$ exploiting the approach presented in Section 3.2.2. We construct a contrastive feedback with InfoNCE (Chen et al., 2020) to provide the positive contrastive signal as defined in Formula 7.
$\mathcal{L}_{c_{1}}=\sum_{i=1}^{B}-\log\frac{\exp(\langle\mathbf{e}^{\text{annotated}}_{i},\mathbf{e}^{\text{rollout}}_{i}\rangle/\tau)\odot\mathcal{M}_{1}}{\sum_{j=1}^{B}\exp(\langle\mathbf{e}^{\text{annotated}}_{i},\mathbf{e}^{\text{rollout}}_{j}\rangle/\tau)},$ (7)
where $\mathcal{M}_{1}$ represents a binary mask, in which each element takes the value 1 if the corresponding CoT leads to a correct answer, and 0 otherwise. The notation $\langle\cdot,\cdot\rangle$ denotes the inner product.
Negative Signal
We devise a scheme to utilize the signal within the negative CoT as well. We denote the annotated CoTs and the associated negative CoTs by $\mathbf{c^{\prime}}_{i}^{\text{annotated}}$ and $\mathbf{c^{\prime}}_{i}^{\text{rollout}}$, respectively. Initially, we calculate the Longest Common Subsequence (LCS) of $\mathbf{c^{\prime}}_{i}^{\text{annotated}}$ and $\mathbf{c^{\prime}}_{i}^{\text{rollout}}$. Subsequently, based on the LCS and the parts of the sequence that exclude the LCS, we construct four embeddings, denoted $\mathbf{e}_{i,\text{LCS}}^{\text{annotated}}$, $\mathbf{e}_{i,\text{exc}}^{\text{annotated}}$, $\mathbf{e}_{i,\text{LCS}}^{\text{rollout}}$, and $\mathbf{e}_{i,\text{exc}}^{\text{rollout}}$, respectively. Then, the negative contrastive signal is formulated in Formula 8.
$\mathcal{L}_{c_{2}}=\sum_{i=1}^{B}-\log\frac{\exp(\langle\mathbf{e}_{i,\text{LCS}}^{\text{rollout}},\mathbf{e}_{i,\text{exc}}^{\text{annotated}}\rangle/\tau)\odot\mathcal{M}_{2}}{\sum_{j=1}^{B}\exp(\langle\mathbf{e}_{i,\text{LCS}}^{\text{rollout}},\mathbf{e}_{j,\text{exc}}^{\text{rollout}}\rangle/\tau)},$ (8)
where $\mathcal{M}_{2}$ represents a binary mask, in which each element takes the value 1 if the corresponding CoT leads to a wrong answer, and 0 otherwise.
Optimization
We optimize the following reinforcement learning loss to learn the policy:
$\mathcal{L}=\mathcal{L}_{\text{RL}}+c\left\{\mathcal{L}_{c_{1}}\,\text{or}\>\mathcal{L}_{c_{2}}\right\}$ (9)
where $c$ balances the relative importance of the PPO and contrastive losses during the reinforced fine-tuning process (see detailed algorithms in Appendix).
3.2.4 Embedding-enhanced Partial Reward
In order to further improve the stability and the performance of the contrastive signal, we propose an embedding-enhanced partial reward method.
ReFT (Luong et al., 2024) assigns a partial reward $r(x,y)=0.1$ to the CoT when it is a negative CoT, from which a numerical answer can be extracted. Unlike the partial reward in ReFT, we introduce a fine-grained partial reward by leveraging our unified CoT Embedding, which provides a tool to measure CoT similarity.
$r(x,y)=\langle\mathbf{e}^{\text{annotated}},\mathbf{e}^{\text{rollout}}\rangle\times 0.1+0.2.$ (10)
The inner product $\langle\mathbf{e}^{\text{annotated}},\mathbf{e}^{\text{rollout}}\rangle$ ranges from 1-1 to $1$, leading to a partial reward range of $[0.1,0.3]$. When the CoTs are dissimilar, the inner product approaches 1-1, resulting in a reward close to $0.1$; when they are similar, the reward approaches $0.3$. This strategy encourages well-behaved CoT generation. By assigning differentiated rewards to negative CoTs, the embedding-enhanced partial reward method further improve the stability of the reinforced fine-tuning process and the final performance of LLMs.
Method SVAMP GSM8K
#Train Samples 3076 7465
#Test Samples 1000 1319
Table 2: Statics of the train and test datasets.
4 Experiments
In this section, we present the experimental results. First, we present our experiment setup. Then, we demonstrate the evaluation of CARFT compared with SFT, ReFT, and Dr.GRPO. Afterward, we present an ablation study.
4.1 Experimental Setup
We conduct experiments on two publicly available datasets: SVAMP (Patel et al., 2021) and GSM8K (Cobbe et al., 2021). Table 2 presents the key statistics of SVAMP and GSM8K. For the reasoning process, we leverage the CoT annotations from (Luong et al., 2024), which were generated based on few-shot prompting (Wei et al., 2022; Gao et al., 2023) with GPT-3.5-turbo (OpenAI, 2023). Our experiments are conducted based on two foundation models: CodeLlama-7B (Rozière et al., 2023) and Qwen2.5-7B-Instruct (Team, 2024). We evaluate CARFT in comparison with three baseline approaches: SFT, ReFT (Luong et al., 2024), and Dr.GRPO (Liu et al., 2025). ReFT is a state-of-the-art RL approach for LLM fine-tuning. As an advanced extension of GRPO (Shao et al., 2024), Dr.GRPO demonstrates excellent performance on R1-like (DeepSeek-AI, 2025) tasks. See setup details in Appendix.
4.2 Evaluation of CARFT
As illustrated in Table 3, CARFT significantly outperforms SFT and ReFT across different models (up to 10.15% on average). With the SVAMP dataset, CARFT yields substantial accuracy enhancements compared with SFT. Precisely, the accuracy escalates from 62.3% to 64.8% and from 86.9% to 88.0%, with absolute increments of 2.5% and 1.1% for CodeLlama and Qwen2.5-Instruct, respectively. Moreover, with the GSM8K dataset, CARFT showcases remarkable improvements as well. The accuracy climbs from 43.82% to 50.95% and from 80.67% to 84.31%, corresponding to absolute boosts of 7.13% and 3.64% for CodeLlama and Qwen2.5-Instruct, respectively.
Table 3 further reveals that ReFT can outperform the SFT baseline once the training process stabilizes. Nevertheless, the performance of ReFT remains inferior to CARFT. In addition, the experimental results demonstrate that ReFT is plagued by the model collapse issue, which significantly undermines its effectiveness. Furthermore, we find that when model collapse occurs, the performance of ReFT lags far behind that of SFT (up to 14.56%) and CARFT (up to 18.2%).
Method Size SVAMP GSM8K Average
CodeLlama + SFT 7B 62.3% 43.82 % 53.06%
CodeLlama + ReFT (Luong et al., 2024) 7B 62.5% 50.27% 56.39%
CodeLlama + CARFT 7B 64.8% 50.95% 57.88%
Qwen2.5-Instruct + SFT 7B 86.9% 80.67% 83.79%
Qwen2.5-Instruct + ReFT (Luong et al., 2024) 7B 85.9% 66.11% 76.01%
Qwen2.5-Instruct + CARFT 7B 88.0% 84.31% 86.16%
Table 3: Evaluation Accuracy of Various Methods on the SVAMP and GSM8K Datasets.
Figure 3: Accuracy curves of various methods on SVAMP dataset and Qwen2.5-7B-Instruct backbone.
Furthermore, Figures 3 depicts the accuracy curves of different approaches on the SVAMP dataset, with Qwen2.5-7B-Instruct serving as the backbone model. These results indicate that ReFT undergoes model collapse after undergoing fine-tuning for just one epoch. In addition, we find that when using Qwen2.5-7B-Instruct, SFT is prone to an unstable tuning process, as evidenced by the decline in its accuracy as the fine-tuning process advances. In contrast, CARFT exhibits remarkable stability and superb performance throughout the entire training process. This outstanding performance can be attributed to our contrastive feedback mechanism as presented in Section 3.2.3, which offers reference signals for the generation of CoTs.
Figure 4: Accuracy curves of various methods on SVAMP dataset and CodeLlama-7B backbone.
As demonstrated in Figure 4, CARFT consistently outperforms ReFT during the fine-tuning process and converges rapidly, swiftly reaching peak accuracy values. Nevertheless, the figure also suggests that CARFT is potentially vulnerable to unstable fine-tuning process. See additional experimental results in Appendix.
As shown in Table 4, CARFT with embedding-enhanced partial reward enabled significantly outperforms all baseline approaches in terms of both accuracy (up to 0.5% compared with Dr.GRPO and 1.7% compared with ReFT) and efficiency (up to 30.62% compared with Dr.GRPO). Interestingly, Dr.GRPO also surpasses ReFT in terms of performance metrics (1.2%), which is accompanied with a considerable increase in computational time. Specifically, Dr.GRPO relies on significant computing resources due to the generation of a larger number of CoTs.
Method Accuracy Time Cost(hours)
ReFT 62.5% 14.12
Dr.GRPO 63.7% 24.49
CARFT 64.2% 16.99
Table 4: Evaluation Accuracy of Various Methods on the SVAMP Datasets, based on CodeLlama-7B.
4.3 Ablation Study
Figure 5: Accuracy of CARFT with positive signal and negative signal, based on the SVAMP dataset and with the CodeLlama-7B as the backbone model.
Figure 6: Accuracy Curve of CARFT with positive signal and negative signal, based on the SVAMP dataset and with the CodeLlama-7B as the backbone model.
Positive Signal versus Negative Signal
We conduct an ablation study to show the impact of positive and negative contrastive signals. As shown in Figure 5, CARFT outperformances ReFT with both positive (2.30% higher) and negative contrastive (0.4% higher) signals. Notably, the positive signal demonstrates a more pronounced performance gain (1.9% higher) compared to its negative counterpart. Due to its excellent performance, we employ the positive signal in our experiments.
Figure 7: Accuracy of CARFT with different $c$, based on the SVAMP dataset and with the CodeLlama-7B as the backbone model.
Robustness
To assess the robustness of the proposed method, we perform an ablation study on the contrastive loss coefficient $c$. As illustrated in Figure 7, by systematically varying the value of $c$ within the range from $5\times 10^{-4}$ to $1.5\times 10^{-3}$, we observe that CARFT consistently outperforms the SFT and ReFT baseline across all tested values. This consistent superiority in performance strongly validates the robustness of CARFT and demonstrates its resilience to changes in the contrastive loss coefficient.
Stability
Figure 8: Accuracy Curve of CARFT Partial Rewards, based on the SVAMP dataset and with the CodeLlama-7B as the backbone model.
To further enhance the stability of the reinforced fine-tuning process, we propose an embedding-enhanced partial reward method, as described in Section 3.2.4. As shown in Figure 8, this approach effectively improves training stability. The tuning process of CARFT using the embedding-enhanced partial reward method achieves a final accuracy that is 0.5% higher than that of the baseline without the method. Moreover, CARFT with this enhancement exhibits a more stable accuracy improvement curve. CARFT with embedding-enhanced partial reward enabled achieves a peak accuracy of 64.2%, which also corresponds to significant improvements over ReFT, with gains of up to 1.7%.
5 Conclusions
In this paper, we propose a novel contrastive learning-based framework with annotated CoTs, i.e., CARFT, to enhance the reasoning capabilities of LLMs. We propose generating contrastive signals from both positive and negative CoTs while incorporating annotated CoTs. In order to further improve the stability of the reinforced fine-tuning process, we propose a novel embedding-enhanced partial reward method. Extensive experimental results demonstrate significant advantages of CARFT in terms of performance (up to 10.15%) and efficiency (up to 30.62%). In addition, CARFT corresponds to better stability during reinforced fine-tuning compared with existing approaches.
Limitations
CARFT requires additional computational overhead to compute the embeddings of CoTs, to achieve excellent. As a result, it consumes a longer computational time compared to the ReFT and SFT. However, CARFT needs less computational time than Dr.GRPO as CARFT needs less on-policy sampled CoTs. In addition, CARFT is designed to exploit a centralized annotated CoT dataset. The annotated datasets may be stored in multiple data centers or devices, which may hinder the application of CARFT with decentralized data.
In this work, we have restricted the context to fewer than 1024 tokens; we plan to explore improving reasoning in long-context scenarios (Zhu et al., 2024).
Appendix A Appendix
A.1 Experimental Settings
All experiments are conducted on an ensemble of 8 H100-80GB GPUs. Given that reinforced fine-tuning for reasoning tasks is inherently time-intensive, we utilize FlashAttention (Dao et al., 2022; Dao, 2024) and DeepSpeed Zero stage 3 (Rajbhandari et al., 2020; Rasley et al., 2020) to expedite the fine-tuning process. These technologies enable us to scale up the batch size, thereby enhancing computational efficiency. Additionally, we utilize the HuggingFace Alignment Handbook (Tunstall et al., ) and the TRL library (von Werra et al., 2020) as methodological guides to streamline the fine-tuning implementation.
To ensure consistency and comparability across experiments, we adopt a structured hyperparameter configuration strategy. During the warmup phase, we initialize training with a batch size of 64 and a learning rate of 1e-5. This learning rate is then adjusted to 3e-7 during the reinforcement fine-tuning stage to stabilize the optimization process. We maintain a batch size of 64 for all models on the SVAMP dataset. On GSM8K, we tailor the batch size to each model’s computational characteristics: 64 for Qwen2.5-7B-Instruct and 96 for CodeLlama-7B, balancing memory efficiency and training throughput. In reinforcement learning components, we set the KL divergence coefficient to 0.05 to regulate policy updates and employ a temperature parameter ($\tau$) of 0.2 and c=1e3c=1e-3 in the contrastive learning loss to control embedding similarity. We set the dimension of the projected embedding to 64. For PPO optimization, we configure $\lambda=1$, $\gamma=0.95$, $\alpha=5$, $\epsilon=0.2$, and $U=2$. For Dr.GRPO, to ensure a fair comparison, we set the parameter $G$ in Dr.GRPO to 2, which matches the maximum number of CoTs in CARFT at each step. We set the reward $r(x,y)$ to 1 if the answer is correct, and 0 otherwise. We also adopt a partial reward scheme (Le et al., 2022), setting the reward to 0.1 in cases where a numerical answer can be extracted but is incorrect.
Training epoch limits are determined based on empirical convergence behavior. For the SFT baseline, we cap training at 60 epochs due to its tendency to be unstable; beyond this point, additional epochs yield diminishing returns. To ensure a fair comparison across methods, for ReFT, CARFT, and Dr.GRPO, we fine-tune the base model for 4 epochs and select the best checkpoint for reinforced fine-tuning. All of these approaches are then trained for 70 epochs, allowing sufficient iterations for convergence while maintaining experimental rigor.
A.2 More Experiments
Batch Size Accuracy
64 51.48%
96 50.95%
Table 5: CARFT with different batch size, based on CodeLlama-7B model and GSM8K dataset.
Batch Size
In our experiments, we utilized FlashAttention (Dao et al., 2022; Dao, 2024) and DeepSpeed Zero stage 3 (Rajbhandari et al., 2020; Rasley et al., 2020) to accelerate the fine-tuning process with a large batch size. To evaluate how batch size affects model performance, we conducted a systematic ablation study. As shown in Table 5, increasing the batch size can degrade the performance of large language models (LLMs). Specifically, enlarging the batch size from 64 to 96 led to a drop in accuracy from 51.48% to 50.95%.
This suggests that reducing the batch size may be a viable strategy for achieving better performance. It is worth emphasizing that all experiments were carried out with consistent batch size configurations to ensure a fair and valid comparison.
Explains of the Accuracy Curve
Figure 9: Accuracy curves of various methods on GSM8K dataset and Qwen2.5-7B-Instruct backbone.
Figure 9 indicates that ReFT also suffers from model collapse, which yields poor results. CARFT shows strong stability across the whole fine-tuning process and outperforms both SFT and ReFT significantly.
Figure 10: Accuracy curves of various methods on GSM8k dataset and CodeLlama-7B backbone.
Figure 10 presents the accuracy of various methods on the GSM8K dataset using the CodeLlama-7B model as the backbone. We observed that further training did not lead to performance improvements in the SFT (Supervised Fine-Tuning) phase, so we terminated the training early. The figure also demonstrates that CARFT outperforms ReFT with a higher convergency accuracy.
Figure 11: Accuracy Curve of CARFT with different $c$, based on the SVAMP dataset and with the CodeLlama-7B as the backbone model.
Figure 11 illustrates the accuracy curves of CARFT under different values of parameter $c$. It can be observed that CARFT attains the optimal performance when c=1e3c=1e-3.
Figure 12: RL loss curve for CARFT with the GSM8K dataset and CodeLlama-7B serving as the backbone model.
Figure 13: Constrastive loss curve for CARFT with the GSM8K dataset and CodeLlama-7B serving as the backbone model.
Figure 14: RL loss curve for ReFT with the GSM8K dataset and CodeLlama-7B serving as the backbone model.
Loss
We present the RL (Reinforcement Learning) and contrastive learning loss curves for CARFT and ReFT models in Figures 12, 13, 14, 15, 16, and 17. These results are based on the GSM8K dataset and utilize the CodeLlama-7B and Qwen2.5-7B-Instruct base models. We make the following observations: (1) As shown in Figure 12 and Figure 14, when using CodeLlama-7B as the base model, CARFT and ReFT exhibit similar loss curves. (2) In contrast, when Qwen2.5-7B-Instruct is used as the base model, the loss curves of CARFT and ReFT differ significantly. In particular, ReFT displays a fluctuating pattern, suggesting instability during fine-tuning. (3) Furthermore, as seen in Figure 13 and Figure 16, the contrastive loss varies across models, indicating differences in the CoT embedding spaces learned by each model.
Figure 15: RL loss curve for CARFT with the GSM8K dataset and Qwen2.5-7B-Instruct serving as the backbone model.
Figure 16: Contrastive loss curve for CARFT with the GSM8K dataset and Qwen2.5-7B-Instruct serving as the backbone model.
Figure 17: RL loss curve for ReFT with the GSM8K dataset and Qwen2.5-7B-Instruct serving as the backbone model.
1
Input : Tuples of (question, CoT, answer): $\mathcal{D}_{\text{train}}=\{(\mathbf{x},\mathbf{c},\mathbf{y}\}$, Number of RL steps: $T$, Number of updates per RL step:$U$, Initial policy: $\mathbf{\pi}_{\theta}^{0}$.
Output : Final Policy: $\mathbf{\pi}_{\theta}$
2 for $i\leftarrow 1$ to $T$ do
 $\mathbf{x},\mathbf{c},\mathbf{y}\sim\mathcal{D}_{\text{train}}$   // Sample training data from $\mathcal{D}_{\text{train}}$
 $\hat{\mathbf{c}}\sim\pi_{\theta}$   // On-policy CoT sampling
3
[…]
4
 $\mathbf{e}^{\text{annotated}}\leftarrow\mathbf{c}$,  $\hat{\mathbf{e}}^{\text{rollout}}\leftarrow\hat{\mathbf{c}}$   // Construct CoT Embeddings
5
6  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{1}$
7 for $i\leftarrow 1$ to $U$ do
    $\theta,\phi\leftarrowOptimization$_step($\mathcal{L}$)   // Equation 9
8
9   end for
10
11 end for
return $\pi_{\theta}$
Algorithm 1 CARFT with Positive Signal
Input : Tuples of (question, CoT, answer): $\mathcal{D}_{\text{train}}=\{(\mathbf{x},\mathbf{c},\mathbf{y}\}$, Number of RL steps: $T$, Number of updates per RL step:$U$, Initial policy: $\mathbf{\pi}_{\theta}^{0}$.
Output : Final Policy: $\mathbf{\pi}_{\theta}$
1 for $i\leftarrow 1$ to $T$ do
 $\mathbf{x},\mathbf{c},\mathbf{y}\sim\mathcal{D}_{\text{train}}$   // Sample training data from $\mathcal{D}_{\text{train}}$
 $\hat{\mathbf{c}}\sim\pi_{\theta}$   // On-policy CoT sampling
[…]
2
 $\mathbf{e}^{\text{annotated}}_{\text{LCS}},\mathbf{e}^{\text{annotated}}_{\text{exc}}\leftarrow\mathbf{c}$,  $\hat{\mathbf{e}}^{\text{rollout}}_{\text{LCS}},\hat{\mathbf{e}}^{\text{rollout}}_{\text{exc}}\leftarrow\hat{\mathbf{c}}$   // Construct CoT Embeddings
3
4  Compute $\sigma_{t},\hat{A}_{t},\hat{R}_{t},\mathcal{M}_{2}$
5 for $i\leftarrow 1$ to $U$ do
    $\theta,\phi\leftarrowOptimization$_step($\mathcal{L}$)   // Equation 9
6
7   end for
8
9 end for
return $\pi_{\theta}$
Algorithm 2 CARFT with Negative Signal
A.3 CARFT with Positive Signal
[…]
A.4 CARFT with Negative Signal
[…]
The masked InfoNCE loss then leverages $\mathbf{e}^{\text{rollout}}_{\text{LCS}}$, $\mathbf{e}^{\text{annotated}}_{\text{exc}}$, and $\hat{\mathbf{e}}^{\text{rollout}}_{\text{exc}}$ to provide feedback for training.